OpenStudy (anonymous):

assume: n is a positve integer, p is a prime number if 2n-1 = p show that n is also prime

OpenStudy (mrhoola):

hmmm . 'n' is not always prime for certain values of p . it works for when p <= 13 .

OpenStudy (kmeis002):

Did you mean \(2^n-1\)? As MrHoola said, a simple counter example is p = 17, n = 9.

OpenStudy (anonymous):

yeah \[2^{n}-1 = p\]

OpenStudy (kmeis002):

Assume \(n \) is not prime, then \(n = ab\) implying that: \[2^n-1 = 2^{ab}-1 = (2^a)^b-1\] We can factor \((2^a)^b-1\) as: \[ p= ((2^a)^b-1) = (2^a-1)((2^a)^{b-1} +(2^a)^{b-2} +...+1)\] Note that this implies that \(p\) is a product of 2 numbers and not prime.